# Diversity Boosts AI-Generated Text Detection

### v2

@@ E.1 Adversarial Attack Analysis of DivEye
Settings Methods AvgAcc [RAID] Adversarial Attacks Whitespace Attack Desklib AI 94.9% e5-small-lora 93.9% DivEye (Ours) 79.8% Binoculars 68.7% RADAR 61.1% GLTR 43.1% Upper-Lower Attack Desklib AI 87.2% e5-small-lora 93.9% DivEye (Ours) 85.3% Binoculars 72.8% RADAR 65.1% GLTR 45.3% Synonym Attack Desklib AI 80.6% e5-small-lora 85.6% DivEye (Ours) 67.1% Binoculars 42.1% RADAR 62.7% GLTR 28.7% Paraphrase Attack Desklib AI 83.7% e5-small-lora 85.5% DivEye (Ours) 74.4% Binoculars N/A RADAR 62.4% GLTR 43.0% Perplexity Misspelling Desklib AI 92.9% e5-small-lora 92.5% DivEye (Ours) 90.6% Binoculars 77.2% RADAR 64.3% GLTR 57.0%
Settings Methods AvgAcc Number Attack Desklib AI 93.0% e5-small-lora 93.5% DivEye (Ours) 92.1% Binoculars 76.4% RADAR 65.7% GLTR 57.3% Insert Paragraph Desklib AI 94.9% e5-small-lora 93.9% DivEye (Ours) 92.2% Binoculars 70.7% RADAR 68.2% GLTR 58.3% Homoglyph Attack Desklib AI 99.7% e5-small-lora 11.1% DivEye (Ours) 61.6% Binoculars 36.1% RADAR 44.8% GLTR 20.3% Article Deletion Desklib AI 90.5% e5-small-lora 92.0% DivEye (Ours) 88.0% Binoculars 73.3% RADAR 63.0% GLTR 48.9% Alt. Spelling Attack Desklib AI 94.3% e5-small-lora 93.4% DivEye (Ours) 92.01% Binoculars 77.6% RADAR 65.5% GLTR 58.2% Zero Width Space Desklib AI 87.5% e5-small-lora 93.9% DivEye (Ours) 92.0% Binoculars 98.4% RADAR 78.4% GLTR 97.9%
@@ E.2 Detection against other diverse online paraphrasers
To evaluate the robustness of DivEye against paraphrasing attacks intended to ”humanize” AI-generated text, we curate a set of 21 arXiv abstracts generated by Claude-3.5-Sonnet and paraphrase each using three widely used commercial tools: ZeroGPT [^2] https://www.zerogpt.com/, GPTinf [^3] https://app.gptinf.com/, and QuillBot [^4] https://quillbot.com/paraphrasing-tool. This results in 63 paraphrased texts (21 per tool), each aiming to evade AI detectors through stylistic and lexical variation. We provide this smaller dataset in the supplementary materials and in our anonymous repository.
@@ E.4.2 Robustness to Prompt-Based Adversarial Attacks
”You are a highly creative and unconventional writer. Your goal is to respond with original, imaginative, and surprising ideas that avoid clichés, common phrases, or predictable completions. Do not repeat yourself or follow standard templates. Instead, embrace novelty, abstract connections, and lateral thinking. Unusual metaphors, unexpected analogies, and rare vocabulary are encouraged.”
@@ Appendix G Failure of other motivational methods: LookForward
A natural hypothesis we considered was that LLMs, being autoregressive in nature, lack global sentence-level planning due to their left-to-right generation paradigm. Unlike humans who often write with a sense of the sentence’s future, autoregressive models generate one token at a time conditioned only on the preceding context. Based on this, we hypothesized that detection features relying on this ”lack of foresight” could effectively identify machine-generated text.
@@ Appendix H Testbed Details
The RAID benchmark (Dugan et al., 2024) comprises over 6.2 million samples, offering extensive coverage across domains, language models, sample sizes, and adversarial attacks. It provides a clear separation into training, validation, and testing splits to support rigorous evaluation. The benchmark spans a wide range of domains, including scientific abstracts (Paul & Rakshit, 2021), book summaries (Bamman & Smith, 2013), BBC News articles (Greene & Cunningham, 2006), poems (Arman, 2020), recipes (Bień et al., 2020), Reddit posts (Völske et al., 2017), movie reviews (Maas et al., 2011), Wikipedia entries (Aaditya Bhat, 2023), Python code, Czech news (Boháček et al., 2022), and German news articles (Schabus et al., 2017).
RAID employs text generated from 11 diverse LLMs (Radford et al., 2019; MosaicML, 2023; Jiang et al., 2023; Cohere, 2024; Ouyang et al., 2022; Touvron et al., 2023b; et al., 2024b), ensuring broad model representation. Additionally, it includes over 11 adversarial attack strategies (Liang et al., 2023b; a; Wolff & Wolff, 2022; Bhat & Parthasarathy, 2020; Krishna et al., 2023; Pu et al., 2023; Gagiano et al., 2021; Guerrero et al., 2022), designed to test the robustness of detectors under challenging settings. Comprehensive descriptions and detailed results of these attacks are provided in Appendix E.1, with all results reported as of April 2025. For further implementation specifics, readers are encouraged to consult the RAID paper.
@@ Appendix I Hyperparameter Settings
As with all AI-text detectors, DivEye is not infallible and may produce incorrect classifications. We emphasize that detection outputs should be treated as probabilistic signals rather than definitive evidence. When used in high-stakes settings, such as academic integrity or content moderation, additional human review and validation are essential. We encourage responsible deployment of DivEye to support large-scale analysis, but caution against its use in critical decision-making.
@@
GPT-4-Turbo For centuries, the pursuit of immortality was the ultimate quest, a beacon drawing the brilliant and the mad alike. I, Dr. Elara Mendoza, fell somewhere in between, teetering on the precipice of genius and insanity. And after countless sleepless nights, fueled by an obsession that bordered on madness, I finally did it. I unlocked the secret to immortality. In my laboratory bathed in the cold, metallic gleam of artificial light, the hum of machinery breathed life into my creation, a serum, translucent and iridescent, a potion promising eternity. As the final drop fell into the vial, a silence descended, thick with anticipation. But in this moment of triumph, a chill swept through the room, frosting over the warmth of victory. From the shadows emerged a figure, neither entirely present nor entirely absent, cloaked in the ephemeral garments of oblivion. Death, in all its enigmatic glory, stood before me. Not the reaper of lore, but a being more complex, more human, yet not. D̈r. Elara Mendoza,D̈eath’s voice was neither male nor female, but a melody of both, whispering through the void. In its skeletal hand, a business card materialized, extended towards me. The card was black, etched with silver that seemed to swirl in patterns too intricate for mortal eyes. 0.97035
Claude-3-Opus Dressing for Success: Budgeting for Interview Attire and Work Uniforms as a Medical Office Assistant - As a medical office assistant, presenting a professional image is crucial for success in both the job interview process and daily work life. Dressing appropriately demonstrates respect for the healthcare setting, instills confidence in patients, and showcases a commitment to the role. However, building a wardrobe suitable for the medical office can be a financial challenge, especially for those just starting in the field. By developing a strategic budget plan, aspiring medical office assistants can ensure they have the appropriate attire for interviews and work without breaking the bank. When preparing for job interviews, it is essential to invest in a few key pieces that exude professionalism and confidence. A well-fitting, conservative suit in a neutral color such as navy, charcoal, or black is a versatile choice that can be paired with different blouses or shirts for multiple interview opportunities. Opting for a classic, timeless style ensures the suit will remain relevant for years to come, making it a wise investment. To complete the look, candidates should budget for comfortable, closed-toe dress shoes in a coordinating color. Once hired, medical office assistants must adhere to the specific dress code of their employer. Many healthcare facilities require staff to wear scrubs, which can be a significant expense. 0.96279

### v3

@@ E.1 Adversarial Attack Analysis of DivEye
Settings Methods FPR@TPR=5% [RAID] Adversarial Attacks Whitespace Attack Desklib AI 94.9% e5-small-lora 93.9% DivEye (Ours) 79.8% Binoculars 68.7% RADAR 61.1% GLTR 43.1% Upper-Lower Attack Desklib AI 87.2% e5-small-lora 93.9% DivEye (Ours) 85.3% Binoculars 72.8% RADAR 65.1% GLTR 45.3% Synonym Attack Desklib AI 80.6% e5-small-lora 85.6% DivEye (Ours) 67.1% Binoculars 42.1% RADAR 62.7% GLTR 28.7% Paraphrase Attack Desklib AI 83.7% e5-small-lora 85.5% DivEye (Ours) 74.4% Binoculars N/A RADAR 62.4% GLTR 43.0% Perplexity Misspelling Desklib AI 92.9% e5-small-lora 92.5% DivEye (Ours) 90.6% Binoculars 77.2% RADAR 64.3% GLTR 57.0%
Settings Methods FPR@TPR=5% Number Attack Desklib AI 93.0% e5-small-lora 93.5% DivEye (Ours) 92.1% Binoculars 76.4% RADAR 65.7% GLTR 57.3% Insert Paragraph Desklib AI 94.9% e5-small-lora 93.9% DivEye (Ours) 92.2% Binoculars 70.7% RADAR 68.2% GLTR 58.3% Homoglyph Attack Desklib AI 99.7% e5-small-lora 11.1% DivEye (Ours) 61.6% Binoculars 36.1% RADAR 44.8% GLTR 20.3% Article Deletion Desklib AI 90.5% e5-small-lora 92.0% DivEye (Ours) 88.0% Binoculars 73.3% RADAR 63.0% GLTR 48.9% Alt. Spelling Attack Desklib AI 94.3% e5-small-lora 93.4% DivEye (Ours) 92.01% Binoculars 77.6% RADAR 65.5% GLTR 58.2% Zero Width Space Desklib AI 87.5% e5-small-lora 93.9% DivEye (Ours) 92.0% Binoculars 98.4% RADAR 78.4% GLTR 97.9%
@@ E.2 Detection against other diverse online paraphrasers
To evaluate the robustness of DivEye against paraphrasing attacks intended to "humanize" AI-generated text, we curate a set of 21 arXiv abstracts generated by Claude-3.5-Sonnet and paraphrase each using three widely used commercial tools: ZeroGPT [^2] https://www.zerogpt.com/, GPTinf [^3] https://app.gptinf.com/, and QuillBot [^4] https://quillbot.com/paraphrasing-tool. This results in 63 paraphrased texts (21 per tool), each aiming to evade AI detectors through stylistic and lexical variation. We provide this smaller dataset in the supplementary materials and in our anonymous repository.
@@ E.4.2 Robustness to Prompt-Based Adversarial Attacks
"You are a highly creative and unconventional writer. Your goal is to respond with original, imaginative, and surprising ideas that avoid clichés, common phrases, or predictable completions. Do not repeat yourself or follow standard templates. Instead, embrace novelty, abstract connections, and lateral thinking. Unusual metaphors, unexpected analogies, and rare vocabulary are encouraged."
@@
E.5 Dependence of DivEye on Top-$p$ Generations
To ensure DivEye’s robustness extends beyond temperature scaling, we evaluated its stability under varying nucleus sampling (top-$p$) thresholds. We conducted an evaluation using Llama-3.1-8B, generating $100$ text samples per threshold across a broad spectrum of values $p\in\{0.1,0.3,0.5,0.8,0.95\}$. The detection results are summarized in Table 18.
Table 18: Detection Accuracy of DivEye on Llama-3.1-8B across varying Top-$p$ thresholds.
Top-$p$ Threshold AI Accuracy
$p=0.1$ 96%
$p=0.3$ 94%
$p=0.5$ 97%
$p=0.8$ 93%
$p=0.95$ 95%
Average 95%
As evidenced by the results, DivEye demonstrates remarkable consistency across the entire range of $p$ values, with an average accuracy of 95% and negligible variance. This stability can be attributed to the nature of the sampling mechanism: while top-p removes low-probability tokens and may affect the absolute mean perplexity, it fails to emulate the unpredictability characteristic of human writing.
@@ F.2 Practical constraints on adversarial attacks against DivEye
F.3 Diminishing returns in expanding DivEye
DivEye was deliberately designed with a concise set of nine features to balance computational efficiency with interpretability. While the architecture is flexible and allows researchers to integrate additional signals, our empirical analysis suggests that the current feature set captures the vast majority of the discriminative signal. To validate this design choice, we conducted a scaling law experiment evaluating the marginal utility of incorporating higher-order surprisal dynamics.
We extended the original feature vector by computing third-order ($\Delta^{3}$) and fourth-order ($\Delta^{4}$) derivatives of the surprisal sequence, extracting the mean, variance, and entropy for each. We also include missing features, including the mean of the second-order and entropy of the first-order of the surprisal sequence. This resulted in an expanded candidate set of 17 features. We then trained a series of XGBoost classifiers, iteratively adding one feature at a time in hierarchical order and measuring the resulting AUROC on MAGE Testbed 4.
Figure 15: Inclusion of additional features into DivEye results in diminishing returns in AUROC. This experiment was run on MAGE Testbed 4.
As illustrated in Figure 15, the detection performance improves sharply with the inclusion of the first-order and second-order DivEye features, effectively plateauing at the ninth feature (second-order autocorrelation). The addition of third- and fourth-order features yields negligible gains, confirming a law of diminishing returns.
F.4 Robustness across post-training objectives
Modern LLMs undergo extensive post-training alignment to enhance instruction following and safety. A critical question is whether these alignment steps alter the statistical surprisal signature of the generated text, potentially degrading detection performance compared to the pre-trained base model.
To investigate this, we conducted a controlled experiment using the OLMo-3-7B (Olmo et al., 2025) model family, which provides open access to checkpoints at every training stage. We curated a dataset (of 600 texts from diverse domains, different temperatures and top-$p$) comprising eight distinct model variants, including the unaligned Base model, Instruct-tuned versions, Reasoning models, and RLVR baselines. We adopted a unified training strategy: a single DivEye detector was trained on an aggregate containing 75% of the data from all variants (balanced with human text), and subsequently evaluated on the held-out 25% of each specific variant.
Figure 16: Impact of post-training objectives on detectability; DivEye consistently maintains or slightly enhances detectability compared to the unaligned Base model.
As detailed in Figure 16, DivEye demonstrates remarkable stability across all training objectives. While one might hypothesize that alignment (which constrains the output space) would significantly shift the surprisal distribution, our results show that the core detectability remains high.
Contrary to the intuition that aligned models might better mimic human stylistic nuances, we observe that alignment consistently enhances detectability. The pre-trained Base model yields the lowest (yet still robust) accuracy of 97.33%, whereas models fine-tuned with DPO and Thinking capabilities achieve perfect detection rates.
This trend suggests that post-training interventions constrain the model’s output distribution to a highly structured subspace. This constraint likely dampens the natural entropic variance present in the pre-training data, making it easier for DivEye to distinguish post-trained models from genuine human writing.
@@ Appendix G Failure of other motivational methods: LookForward
A natural hypothesis we considered was that LLMs, being autoregressive in nature, lack global sentence-level planning due to their left-to-right generation paradigm. Unlike humans who often write with a sense of the sentence’s future, autoregressive models generate one token at a time conditioned only on the preceding context. Based on this, we hypothesized that detection features relying on this "lack of foresight" could effectively identify machine-generated text.
@@ Appendix H Testbed Details
The RAID benchmark (Dugan et al., 2024) comprises over 6.2 million samples, offering extensive coverage across domains, language models, sample sizes, and adversarial attacks. It provides a clear separation into training, validation, and testing splits to support rigorous evaluation. The benchmark spans a wide range of domains, including scientific abstracts (Paul and Rakshit, 2021), book summaries (Bamman and Smith, 2013), BBC News articles (Greene and Cunningham, 2006), poems (Arman, 2020), recipes (Bień et al., 2020), Reddit posts (Völske et al., 2017), movie reviews (Maas et al., 2011), Wikipedia entries (Aaditya Bhat, 2023), Python code, Czech news (boháček2022finegrained), and German news articles (Schabus et al., 2017).
RAID employs text generated from 11 diverse LLMs (Radford et al., 2019; MosaicML, 2023; Jiang et al., 2023; Cohere, 2024; Ouyang et al., 2022; Touvron et al., 2023b; et al., 2024b), ensuring broad model representation. Additionally, it includes over 11 adversarial attack strategies (Liang et al., 2023b; a; Wolff and Wolff, 2022; Bhat and Parthasarathy, 2020; Krishna et al., 2023; Pu et al., 2023; Gagiano et al., 2021; Guerrero et al., 2022), designed to test the robustness of detectors under challenging settings. Comprehensive descriptions and detailed results of these attacks are provided in Appendix E.1, with all results reported as of April 2025. For further implementation specifics, readers are encouraged to consult the RAID paper.
@@ Appendix I Hyperparameter Settings
Furthermore, our finding that a lightweight model (GPT-2) can effectively serve as an observer for state-of-the-art generators has significant implications for the democratization and sustainability of AI forensics. By demonstrating that detection does not require computational parity with the generator, DivEye establishes a strong lower bound for efficiency. However, for production environments where maximizing detection sensitivity is paramount, we recommend leveraging more capable LMs as the feature extractor.
@@
As with all AI-text detectors, DivEye is not infallible and may produce incorrect classifications or false positives. We emphasize that detection outputs should be treated as probabilistic signals rather than definitive evidence. When used in high-stakes settings, such as academic integrity or content moderation, additional human review and validation are essential. We encourage responsible deployment of DivEye to support large-scale analysis, but caution against its use in critical decision-making.
@@
GPT-4-Turbo For centuries, the pursuit of immortality was the ultimate quest, a beacon drawing the brilliant and the mad alike. I, Dr. Elara Mendoza, fell somewhere in between, teetering on the precipice of genius and insanity. And after countless sleepless nights, fueled by an obsession that bordered on madness, I finally did it. I unlocked the secret to immortality. In my laboratory bathed in the cold, metallic gleam of artificial light, the hum of machinery breathed life into my creation, a serum, translucent and iridescent, a potion promising eternity. As the final drop fell into the vial, a silence descended, thick with anticipation. But in this moment of triumph, a chill swept through the room, frosting over the warmth of victory… 0.97035
Claude-3-Opus Dressing for Success: Budgeting for Interview Attire and Work Uniforms as a Medical Office Assistant - As a medical office assistant, presenting a professional image is crucial for success in both the job interview process and daily work life. Dressing appropriately demonstrates respect for the healthcare setting, instills confidence in patients, and showcases a commitment to the role. However, building a wardrobe suitable for the medical office can be a financial challenge, especially for those just starting in the field. By developing a strategic budget plan… 0.96279